solve the following question by using bisection method.

correct to the four decimal places.


x² ⁻ 3 = 0


1
Expert's answer
2021-04-19T15:40:19-0400

x2-3=0

Our function becomes f(x) = x2-3 ,to get interval substitute various values as follow

When x = 0 then f(0) =02-3 = -3

When x =1 then f(1) = 12-3 = -2

When x = 2 then f(2) = 22-3 =1

The result changes from Negative value (-2) to positive value (1) hence our interval is (1,2) where a =1 and b=2

Using bisection method,x value is obtained as

x=(a+b)÷2 where a is the lower limit of our interval and b is the upper limit of our interval

We can Perform iterations in the following ways:

x1=(1+2)÷2 =1.5

We know that f(x)=x2-3 ,when x=1.5,then we have

f(1.5)=(1.5)2-3=-0.75

The interval changes based on value of f(x) such the the lower limit of interval (a,b) is replaced by (a+b)÷2 when the value of f(x) is negative and upper limit b of our interval is replaced by (a+b)÷2 when the value of f(x) is positive

Since (-0.75) is a negative value,we replace lower limit of interval (1,2) with x1 value 1.5  to obtain new interval as (1.5,2) to get x2 we use new interval (1.5,2)

x2=(1.5+2)÷2 =1.75

We know that f(x)=x2-3

When x=1.75 then we have

f(1.75)=(1.75)2-3=0.0625 since 0.0625 is a positive values we replace upper limit of interval (1.5,2) with x2 value 1.75 to obtain new interval as (1.5,1.75) and to get x3 we use our new interval (1.5,1.75)

x3=(1.5+1.75)÷2=1.625

We know that f(x)=x2-3

When x=1.625 then we have

f(1.625)=(1.625)2-3=-0.359375 since -0.359375 is a negative value,then we replace the lower limit of interval (1.5,1.75) when our new x value 1.625 to obtain new interval as (1.625,1.75)

To get x4 we use our new interval (1.625,1.75)

x4=(1.625+1.75)÷2=1.6875

We know that f(x)=x2-3

When x=1.6875

f(1.6875)=(1.6875)2-3=-0.15234375 since -0.15234375 is a negative value,then we replace lower limit of our interval (1.625,1.75) with our new x value 1.6875 to obtain new interval as (1.6875,1.75) to get x5 we use our new interval (1.6875,1.75)

x5=(1.6875+1.75)÷2=1.71875

We know that f(x)=x2-3

When x=1.71875

f(1.71875)=(1.71875)2-3=-0.0458984375 since -0.0458984375 is a negative value,then we replace lower limit of interval (1.6875,1.75) with our new x value 1.71875 to obtain new interval as (1.71875,1.75) to get x6 we use our new interval (1.71875,1.75)

x6=(1.71875+1.75)÷2=1.734375

We know that f(x)=x2-3

When x=1.734375

f(1.734375)=(1.734375)2-3=0.008056640625 since 0.008056640625 is a positive value,then we replace upper limit of interval (1.71875,1.75) with our new x value 1.734375 to obtain new interval as (1.71875,1.734375) to get x7 we use our new interval (1.71875,1.734375)

x7=(1.71875+1.734375)÷2=1.7265625

We know that f(x)=x2-3

When x=1.7265625

f(1.7265625)=(1.7265625)2-3=-0.01898193359375 since -0.01898193359375 then we replace lower limit of interval (1.71875,1.734375) with new x value 1.7265625 to obtain new interval as (1.7265625,1.734375) to get x8 we use new interval (1.7265625,1.734375)

x8=(1.7265625+1.734375)÷2=1.73046875

We know that f(x)=x2-3

When x=1.73046875

f(1.73046875)=(1.73046875)2-3=-0.00547790527344 since -0.00547790527344 is a negative value then we replace lower limit of interval (1.7265625,1.734375)with our new x value 1.73046875 to obtain new interval as (1.73046875,1.734375) to get x9 we use new interval (1.73046875,1.734375)

x9=(1.73046875+1.734375)÷2=1.732421875 we know that f(x)=x2-3

When x=1.732421875

f(1.732421875)=(1.732421875)2-3=0.00128555297852 since 0.00128555297852 is a positive value,then we replace upper limit of interval (1.73046875,1.734375) with our new x value 1.732421875 to obtain new interval as (1.73046875,1.732421875) to get x10 we use new interval (1.73046875,1.732421875)

x10=(1.73046875+1.732421875)÷2=1.7314453125 we know that f(x)=x2-3

When x=1.7314453125

f(1.7314453125)=(1.7314453125)2-3=-0.00209712982178 since -0.00209712982178 is a negative value,then we replace lower limit of interval (1.73046875,1.732421875)with new x value 1.7314453125 to obtain new interval as (1.7314453125,1.732421875) to get x11 we use new interval (1.73144553125,1.732421875)

x11=(1.7314453125+1.732421875)÷2=1.73193359375 we know that f(x)=x2-3

When x=1.73193359375 then

f(1.73193359375)=(1.73193359375)2-3=-0.00040602684021 since -0.00040602684021 is a negative value,then we replace lower limit of interval (1.7314453125,1.732421875) with our new x value 1.73193359375 to obtain new interval as (1.73193359375,1.732421875) to get x12 we use new interval (1.73193359375,1.732421875)

x12=(1.73193359375,1.732421875)÷2=1.732177734375 we know that f(x)=x2-3

When x=1.732177734375

f(1.732177734375)=(1.732177734375)2-3=0.00043970346451 since 0.00043970346451 is a positive value,then we replace upper limit of interval (1.73193359375,1.732421875) with our new x value 1.732177734375 to obtain new interval as (1.73193359375,1.732177734375)

To get x13 we use new interval (1.73193359375,1.732177734375)

x13=(1.73193359375,1.732177734375)÷2=1.7320556640625 we know that f(x)=x2-3 ,when x=1.7320556640625,then

f(1.7320556640625)=(1.7320556640625)2-3=0.00006823410987854 since 0.00006823410987854 is a positive value,then we replace upper limit of interval (1.73193359375,1.732177734375) with our new x value 1.7320556640625 to obtain new interval as (1.73193359375,1.7320556640625) to get x14 we use new interval (1.73193359375,1.7320556640625)

x14=(1.73193359375,1.7320556640625)÷2=1.73199462871875

therefore x = 1.7320 correct to four decimal places